Subject: APO 3.5m users committee minutes, Sept 13, 2010 meeting

From: Suzanne Hawley

Submitted: Mon, 20 Sep 2010 13:14:58 -0700 (PDT)

Message number: 1199 (previous: 1198, next: 1200 up: Index)

             APO 3.5-m Users Committee Phonecon, 9/13/10

Attending:  Suzanne Hawley, Mark Klaene, Ed Turner, Bill Ketzeback,
Jon Holtzman, Remy Indebetouw, Scott Anderson, Russet McMillan

Absent:  Al Harper, John Bally, UVa


User feedback, comments from institutional representatives:

Chicago (Al Harper) - no report/not present.

New Mexico State (Jon Holtzman) -  Jon plans to bring a class 
to APO for training in November.

Washington (Scott Anderson) - Scott thanked the APO staff for
helping with the UW class that visited APO during July for training.

Colorado (John Bally) - no report/not present

Princeton (Ed Turner for Michael Strauss) - Ed mentioned that there are 
a group of postdocs that are now using the 3.5m a lot and are also interested
in instrumentation.

Johns Hopkins - no report, Sean Moran just named as new users
committee representative.

Virginia (Remy Indebetouw) - Remy mentioned that a department 
meeting will be held soon and they will be discussing the future including 
the 3.5m.


Discussion of telescope/instruments report - Klaene:

Mark highlighted the recent shutdown activity including a repair to 
the tertiary mirror cell wall.  Even with this we accomplished all 
the items on the shutdown task list and returned to science 2 days early. 
It also appears that we are slowly coming out of the monsoon weather pattern.

 	3.5-m Telescope, Instruments, and CIF Projects Highlights, 6/11/10 through 9/05/10

0) Overview

The monsoons came a little early this year along with the passage nearby of 1 hurricane and 1 tropical storm remnant. This caused a lot of lost time at the end of June due to weather.  The scheduled time in July prior to the start of the shutdown was moderately successful in observing around the monsoonal moisture. We hosted a UW student group and a Planetary Nebula working group here during July.  The summer shutdown was completed a couple days ahead of time and some science programs were given unscheduled time.  Weather in August continued to be inconsistent with clouds,rain and humidity causing moderate loss in time but many partial nights were observable.

1) Telescope

No problems prior to the shutdown. Shutdown activities included M1 washing, M2 and M3 re-aluminizing, M2 and M3 actuator servicing, PMSS servo cleanup and tuning, modification to M3 clamp, and Altitude axis controller software update. There have been very few problems since the shutdown.  The most noticeable is an over-current on the Agile rotator caused by umbilical snags.  This problem has not been completely corrected yet.

A minor TUI upgrade was also released.

2) Instruments

DIS, TripleSpec, NIC-FPS, the Echelle, and Agile ran with no problems.

SPICam is running about 20 deg warmer during the daytime and about 10 deg warmer than usual at night. Currently being investigated.

NIC-FPS vacuum was serviced and had to include ION pump replacement. During this time the Leach controller was at CU to investigate the read noise issue and the spare controller abnormality.  Upon return to the mountain, a couple of weeks of fine tuning was required to get the noise and gain values back in spec.

TSpec vacuum was also serviced during this time with no other changes.

Agile computer was upgraded.

Echelle slitviewer computer was upgraded.

Significant progress was made in completing the private network which enhances our computer security.

3) CIF projects

We will start defining FY 2011 CIF projects in the coming month.  The plan
is to continue to set aside funds for instrument upgrades/replacement.

4) Nick Ule, an NMSU graduate student, is currently being trained as a new Observing Specialist.  He will take the place of Jeff Coughlin, and will be working 1-2 nights/month.


Summer shutdown report and early return to science - Klaene, Hawley

Suzanne conveyed her thanks to all those involved (especially Bill, 
Nick, Joe, Fritz, Ed, and Ben).


Echelle upgrade update - Holtzman

After reviewing the white paper, Jon has focused on three areas: 
the detector, the optics coatings, and the grating. 
The most compelling case for a detector replacement will be if we can 
get one with high QE.  There isn't a big improvement to be gained purely 
on the basis of improved read noise.  He is making inquiries about improved 
QE detectors. There does not appear to be a significant gain from 
changing/adding optical coatings especially given the possible difficult
alignment issues. He is also making inquiries about a replacement grating. 
Suzanne requested a short statement with cost estimates by the next meeting
to be discussed by the Users Committee and presented at the 
ARC Board meeting in November.


Scheduling issues - Hawley, McMillan

There are two issues that have come up recently.  One is a large increase
in the number of partial night programs, and the other, which is related,
is the issue of morning calibrations.  Users should be aware that when
they are scheduled next to a partial night program, they are often getting
additional time compared to what was forwarded by their institutional
scheduler.  Sometimes having time spread out over more half nights is
beneficial -- we try to assess that from the proposals, but users 
should feel free to mention it in their proposals to help us with
scheduling.  As far as morning cals, the setup message has recently
been updated to note that cals that run past 7:30am local (mountain)
time must be requested and approved ahead of time.  The Observing 
Specialists work very long nights during the winter months and staying
past 7:30 is a significant burden.  Also the day crew starts work in 
the dome at that time.  An alternative may be to take cals during the
afternoon before the observing.


Science highlights for annual report - Hawley

Suzanne made the first call for science highlights for the annual report,
which are due by Friday, October 15.  Highlights should include text and 
an accompanying image.  These will be presented in a powerpoint 
presentation so if they can be sent as a jpeg or powerpoint slide 
that reduces our editing.  Highlights should be sent to the institutional 
users committee representative for review, and then forwarded to Suzanne 
and Mark.


NIC-FPS read noise - Hawley

Suzanne mentioned that the current proposal to continue working on the 
NIC-FPS noise issue carries with it some risk and likely places the 
instrument off line for 2-3 months.  Prior to the 16 chl readout upgrade, 
read noise was around 14 electrons with Fowler sampling (NFS) 1.  Now, after 
the upgrade, the read noise is 72 electrons for NFS 1 and can be reduced to 21 
electrons with NFS 8 which is the usual operating mode.  We would like to 
hear from the users if the current detector performance is sufficient for 
their science prior to committing further resources and taking the 
instrument out of service for an extended period. See action item below.


TripleSpec acceptance - Hawley

We plan to formerly accept TripleSpec at the November board meeting. 
The remaining items are minor (spares, temp sensor replacement and minor 
documentation issues) and will be addressed soon.  The instrument itself 
has been working quite well and the documentation is quite good. There is 
a desire by some for software that would provide a 2D wavelength solution, 
however that is outside the scope of the instrument acceptance. Ed Turner 
stated he is in favor of accepting the instrument.


APO futures committee -  Hawley

A call for pre-proposals has recently been issued soliciting ideas for
using the 2.5m and 3.5m telescopes after SDSS III is completed in mid-2014.
The idea is to start an early discussion on the best way to utilize 
the 2.5m telescope while at the same time looking at improvements to 
the 3.5m.  It is important to remember that the cost to operate the 
3.5m is impacted by the 2.5m operations.


other business - group

Suzanne stated that she was recently approached by Phil Muirhead (Cornell) 
as to whether or not we would be interested in an interferometer for 
TripleSpec. Phil has developed an interferometer for the Palomar 
TripleSpec instrument and we will be interested in seeing how it works



[open from previous meeting]:

===> ACTION:  Users Committee members should poll their users to find out 
how important it is to their science programs to track targets within
5 degrees of zenith.

+++> RESPONSE: General consensus was that users would like to go as
high as possible, but that a limit of 85 deg was reasonable.  Mark stated 
that 85 deg will be our target for future upgrades. Currently we start 
seeing moderate to severe effects at 82 deg.

---> DISPOSITION: Closed


===> ACTION: Users Committee members to poll their users to see whether
the current performance of NIC-FPS is adequate for their science.


Next meeting:  The next Users Committee phonecon will be on Monday, 
11 October at 8:30 AM Pacific Time.  The agenda and other materials 
will be sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1199 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to